Determine what is asked. Show all pertinent parts of the solution (GUFSA), make sure to express your answer with correct number of significant figures.


  1. Displacement at t = 3s to t = 5s
  2. Velocity at t = 6s, considering that the initial velocity is 4m/s
1
Expert's answer
2021-10-04T10:05:57-0400

Since there is no equation given, assume the acceleration is 5 m/s2.

1. The displacement at 3 and 5 s (starting at 4 m/s):


"d_3=v_0t_1+\frac{at_1^2}2,\\\space\\
d_5=v_0t_2+\frac{at_2^2}2,\\\space\\
\Delta d=d_5-d_3=77\text{ m}."

The velocity is


"v=v_0+at=34\text{ m/s}."